OpenStudy (anonymous):

Write the standard equation of the circle with center (2, 3) and a diameter of 12

OpenStudy (anonymous):

A (x - 2)2 + (y - 3)2 = 36 B (x + 2)2 + (y + 3)2 = 12 C (x - 2)2 + (y - 3)2 = 6 D (x - 3)2 + (y - 2)2 = 36

OpenStudy (anonymous):

Im thinking its A im not sure tho?

OpenStudy (anonymous):

the diameter is twice the radius, so use the normal equation from the previous questions...with r=6

OpenStudy (anonymous):

Yes so 6 squared would be 36?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so then it would bbe A right? (:

OpenStudy (anonymous):

have confidence! :)

OpenStudy (anonymous):

im scared if i fail i wont graduate :(

OpenStudy (anonymous):

you know this stuff...you're getting them right. :)

OpenStudy (anonymous):

:) THank you

OpenStudy (anonymous):

my pleasure....glad i can help :)